TR= 75Q-4Q^2




Find AR and MR function?




At what level of output and price MR is zero?

1
Expert's answer
2022-03-21T18:47:20-0400

"TR= 75Q- 4Q^2"

"AR= \frac{TR}{Q}= \frac{75Q- 4Q^2}{Q}"

= 75- 4Q

MR= "\frac{\delta TR}{\delta Q}= 75-8Q"

MR= 0

75-8Q= 0

Q= 9.375

P= "\frac{TR}{Q}=\frac{ 75(9.375)- 4(9.375^2)}{9.375}"

= "\frac{351.5625}{9.375}= 37.5"